OpenStudy (anonymous):

the perimeter of a certain square is seven times the length of a side, decreased by 12 find the length of a side

OpenStudy (radar):

perimeter=seven times (side-12) or letting x = side then: 4x=7(x-12). subtracting 4x from both sides and adding 84 to both sides results in: 7x-4x=84 simplifying 3x=84, solving for side x = 28. Proof: 4*28=7*(28-12) 112=112